How would you solve this problem what method would you use? An airplane flies directly overhead at 600 mph. 1 minute later it is seen at an angel of elevation of 35 degrees.
1. How far did the airplane fly during that time?
2. What is the elevation?

Answer by stanbon(75887)

How would you solve this problem what method would you use? An airplane flies directly overhead at 600 mph. 1 minute later it is seen at an angel of elevation of 35 degrees.
---
distance = rate*time = 600(mi/hr)(1/60)hr = 10 miles
---
Draw a right triangle with base = 10, altitude = "h" and an
angle of 35 degrees opposite the altitude.
-------
1. How far did the airplane fly during that time?
figured above: 10 miles
--------------------------------
2. What is the elevation?
tan(35) = h/10
h = 10tan(35)
h = 7.00 miles
----
That's about 7(5280)= 36971 ft.
====================================
Cheers,
Stan H.